OpenStudy (anonymous):

When given the equations of x+4 =y and y-x = -3...would these be considered to be parallel?

OpenStudy (anonymous):

set both equal to y y=x+4 y=x-3 Are the coefficients of x in each equation equal? Let's check. y=(1)x+4 and y=(1)x-3, yep, parallel

OpenStudy (katrinakaif):

Parallel equations have to have the same slope. Convert the equations to y = mx + b format First one already is. The second one - add x to the other side. Since the slope (coeffiienct on x) is both 1 theiy are aparralel.

OpenStudy (anonymous):

yes because cooefficients of x are seen to be same when written in form y = mx + c